OpenStudy (anonymous):

Write the equation of the line, in standard form, that passes through the origin and is parallel to x + y = 6. Include your work in your final answer. I understand the answer is x + y = 0 I want to understand how that is determined. Using point slope form and all that.

OpenStudy (campbell_st):

well it needs an x intercept of 0 and y intercept of 0 and slope of -1 so if you take the given line x + y = 6 then y = -x + 6 so the slope if m = -1 the point is the origin (0,0) then y - 0 = -1(x - 0) just simplify

OpenStudy (anonymous):

So you determine the slope by the points?

OpenStudy (campbell_st):

the slope comes from the original line its in standard form x + y = 6 re=writing in slope intercept form y = -x + 6

OpenStudy (campbell_st):

so the slope is m = -1

OpenStudy (anonymous):

Okay so the -x is the slope? What about the 6? Does it not do anything?

OpenStudy (campbell_st):

no the slope is -1 the general slope intercept form is y = mx + b m is the slope so in the given line m = -1 b is the y-intercept

OpenStudy (anonymous):

Oh I see what I was doing. Thank you. I understand.